# TreeDGS: Aerial Gaussian Splatting for Distant DBH Measurement

### v2

@@ 5 Conclusion
TreeDGS has important limitations. First, it relies on trunk visibility: the breast-height band must be observed in $\sim$2–3+ informative views for SfM/MVS and Gaussian optimization to constrain the stem surface. When this band is fully occluded in most views, the method cannot recover the missing geometry, regardless of how densely the Gaussian field is sampled; sampling density cannot compensate for geometry that is never sufficiently observed in the input imagery. Second, DBH estimation is sensitive to 3D semantic segmentation quality. Because diameter is fit from trunk-isolated points, false-positive trunk labels from nearby vegetation can bias the estimate. Although multi-view opacity weighting mitigates this effect, the pipeline remains subject to the failure modes of the trunk classifier [48]. This sensitivity can also contribute to poorer LiDAR baseline performance in challenging plots: breast-height stem returns with a small number of vegetation false positives can disproportionately bias cylinder fits.
@@
A promising direction is to reduce dependence on post-processing (sampling and segmentation) by incorporating explicit trunk primitives into Gaussian optimization. Jointly optimizing a small set of parameterized stem elements (e.g., tapered cylinders) together with the Gaussian field would enable DBH to be inferred directly from model parameters and would provide a stronger inductive bias under partial trunk visibility. Such a hybrid representation may also support end-to-end supervision when ground-truth field DBH is available for future learning-based systems.

### v3

@@ 5 Discussion
TreeDGS has important limitations. First, it relies on trunk visibility: the breast-height band must be observed in at least a few frames (in practice, $\sim$2+ informative views for SfM/MVS and Gaussian optimization to constrain the stem surface (Figure 17a). When this band is fully occluded in most views, the method cannot recover the missing geometry, regardless of how densely the Gaussian field is sampled; sampling density cannot compensate for geometry that is never sufficiently observed in the input imagery.
Figure 17: Examples of successful and failed DBH fits at breast height. (ac) Failure cases: (a) no unoccluded breast-height views, (b) strongly curved/leaning stem causing biased horizontal slicing, and (c) instance segmentation that merges a target stem with a nearby sapling. (df) Successful cases with clear breast-height visibility and accurate per-slice fits. Yellow circles show the fitted cross-sections on representative slices. Green dots show the segmented trunk points; the estimated DBH (Est.) is colored red for failure cases and blue for successful cases.
Second, DBH estimation is sensitive to 3D semantic segmentation quality. Because diameter is fit from trunk-isolated points, false-positive trunk labels from nearby vegetation can bias the estimate (Figure 17c). Although multi-view opacity weighting mitigates this effect, the pipeline remains subject to the failure modes of the trunk classifier [50]. This sensitivity can also contribute to poorer LiDAR baseline performance in challenging plots: when breast-height stem returns a small number of vegetation false positives can disproportionately bias cylinder fits.
Third, our current DBH estimator makes a morphological assumption that is well satisfied in the managed Pinus taeda plantation used here but may not hold in more complex forests. Specifically, we assume that (i) the lower stem cross-section can be reasonably approximated by a circle (per-slice solid-circle RANSAC), (ii) horizontal slicing is a good proxy for slicing orthogonal to the stem axis (i.e., stems are approximately vertical/straight near breast height), and (iii) diameter decreases monotonically with height over the fitted lower-stem segment (linear taper with a negative-slope prior). In natural or mixed-species forests, stems can be leaning, curved (Figure 17b), buttressed, fluted, forked, or multi-stem, and these cases can violate the above assumptions and introduce bias. Importantly, this limitation is in the geometric measurement model rather than in TreeDGS sampling itself: the opacity-weighted sampling yields a denser, reliability-weighted point set that could support more flexible trunk models. We view extending the measurement stage to such morphologies and validating on mixed-species natural forests as key future work.
Future Work.
@@
A promising direction is to reduce dependence on post-processing (sampling and segmentation) by incorporating explicit trunk primitives into Gaussian optimization. Jointly optimizing a small set of parameterized stem elements (e.g., tapered cylinders) together with the Gaussian field would enable DBH to be inferred directly from model parameters and would provide a stronger inductive bias under partial trunk visibility. Such a hybrid representation may also support end-to-end supervision when ground-truth field DBH is available.
6 Conclusions
We introduced TreeDGS, a UAV RGB-only pipeline that repurposes 3D Gaussian splatting as an opacity-weighted density sampling method for trunk measurement, rather than only a view-synthesis renderer. Building on SfM/MVS initialization, TreeDGS optimizes a Gaussian field and then converts it into a measurement-ready stem point set via depth-aware cumulative-opacity sampling and multi-view reliability weighting, enabling robust circle fitting at breast height. In doing so, our approach bridges a key gap left by prior UAV LiDAR and UAV-SfM DBH pipelines that operate on inherently sparse or view-limited point clouds, and demonstrates that Gaussian splats can support accurate trunk-level metrics from commodity aerial imagery.
Acknowledgments
We thank PotlatchDeltic Corporation for supporting this research by allowing us to conduct data collection and validation on their managed forest stands. We are grateful to Jacob Strunk, Kit Hart, Nathaniel Naumann, Paurava Thakore, Tim Sydor, and Bill Driscoll for coordinating logistics, sharing field context, and assisting with site access and study planning.
Data Availability
The original contributions presented in this study are included in the article. Further inquiries can be directed to the corresponding author.
Conflict of Interest
Authors Belal Shaheen, Minh-Hieu Nguyen, Bach-Thuan Bui, Shubham, Tim Wu, Michael Fairley, Matthew Zane, and Michael Wu were employed by Coolant. James Tompkin is a scientific advisor to Coolant. The authors declare that these affiliations did not influence the study design, data collection, analysis, interpretation of results, or the decision to publish.